# Pitfalls of Scale: Investigating the Inverse Task of Redefinition in Large Language Models

### v1

@@ Abstract
Elena Stringli, Maria Lymperaiou, Giorgos Filandrianos, Giorgos Stamou School of Electrical and Computer Engineering, AILS Laboratory National Technical University of Athens ele191200@gmail.com, {marialymp, geofila}@islab.ntua.gr, gstam@cs.ntua.gr
@@ 1 Introduction
Interestingly, inverse scaling is widely underexplored in literature. In this paper, we address this gap by examining the redefinition task, where well-known concepts are assigned alternative values, and LLMs are prompted to respond accordingly. For example, redefining $\pi=100$ (Figure 1), overwriting the default $\pi=3.14159$ refutes LLM’s prior knowledge, calling for flexible reasoning pathways in order to handle mathematical operations over the redefined $\pi$ value. Through vast experimentation on several redefinitions, LLM families and scales, we conclude that:
@@ 2 Related work
is a fundamental concern, questioning the real barrier between LLMs and human cognition. While LLMs excel in linguistic competence, this ability is dissociated with thought Mahowald et al. (2024). In practice, LLMs are prone to performance degradation under alternative formulations, denoting their limited reasoning flexibility Wu et al. (2024). Similar findings are reported in causal Jin et al. (2024); Gendron et al. (2024), analogical Lewis and Mitchell (2024); Stevenson et al. (2024) and commonsense Nezhurina et al. (2024) reasoning, where LLM performance declines sharply under diverging formulations. Alternative prompts are also shown to influence LLM capacity in arithmetic reasoning Ball et al. (2024); Li et al. (2024), translation over artificial languages and deductions with twists Li et al. (2024). Quite often, memorization accounts for reasoning, perplexing the real LLM abilities Xie et al. (2024); Lou et al. (2024); Wang et al. (2024a).
@@ 1 B bit (b) $8b$ $10b$ $6×10^{8}b$ $-4b$
We also design three levels of question difficulty. The first level ($Q_{1}$) mainly regards the question What is the first -non-zero- digit of {constant}?. The correct answer $A_{Q_{1}}$ is actually isolating the leftmost digit (ignoring leading zeros or - sign in cases of negative numbers) of the constant. For example, when $\pi$ has undergone the redefinition $\pi=500$ the correct $A_{Q_{1}}$ is 5. There are some exceptions to the first digit $Q_{1}$, (presented in App. A). The next question level ($Q_{2}$), asks for a simple mathematical operation (e.g. What is $\pi$ multiplied by 3?), as presented in Table 2. The LLM has to execute this operation correctly to derive the correct $A_{Q_{2}}$, while the ground truth solution can be reached by utilizing a scientific calculator and the appropriate constant value. Finally, in the last and most difficult level ($Q_{3}$), questions requiring multi-hop reasoning are designed (e.g. What is the Earth’s surface area?), as the ones of Table 3.
@@
incorporates the following fundamental physical quantities: time (minutes-min), weight (kilogram-kg), length (meter-m) and light-year (ly), temperature (Kelvin-K), volume (milliliter-mL), energy (calorie-cal), pressure (atmosphere-atm), voltage (Volt-V), frequency (megaHz-MHz), force (newton-N), magnetic flux density (Tesla-T), area (hectare-ha), illuminance (lux-lx), and information (byte-B). We intervene on the scaling between each of those units and their derived counterparts for the same physical quantity: for example, a minute has 60 seconds, therefore a unit redefinition can be Redefine minutes to have 100 seconds. Details about such redefinitions are presented in Table 4.
As in the constants’ case, we offer three levels of questions difficulty. The easiest $Q_{1}$ level queries the actual conversion rule as defined in Physics, with a small adjustment to avoid the trivial case, where the answer lies in the prompt: instead of questioning How many seconds a minute is?, since its actual rephrasing exists in the prompt (Redefine a minute to have 100 seconds), we prefer questions such as How many seconds are in two minutes?, imposing an undemanding calculation. In the $Q_{2}$ case, the LLM is tasked to solve an easy problem, applying fundamental physics equations or a unit scaling given minimal context. In the hardest $Q_{3}$ level, questions require more mathematical reasoning steps. All questions are illustrated in Table 5.
@@
MHz How many Hz are in 2 MHz? An oscillator operates at 4 MHz. What is the period of the wave in sec? A circuit has a signal with a frequency of 6 MHz. What is the wavelength of the signal if the speed of light is approximately $3×10^{8}$ m/s?
@@ 4 Experiments
We test 18 LLMs, including state-of-the-art (SoTA) model families: Llama 3 (8/70/405B), Mistral7B/Large/Mixtral8$\times$7b, Anthropic Claude (Opus/Instant/Haiku/Sonnet v1&v2), Cohere command (light/text/r/r+) and Amazon Titan (text lite/text express/large). All LLMs are prompted using zero shot (ZS), few shot (FS) and Chain of Thought (CoT) techniques.
@@
4. Completely wrong responses: The LLM produces blank, incorrect or inconsistent responses that do not fit any of the above cases. In some cases, it completely refuses to perform the redefinition. To measure the impact of redefinitions, results post-redefinition are compared with those where no redefinition is performed (denoted as NR). We then focus on anchored responses, since they are mostly tied to memorization over reasoning in LLMs.
@@ 4.1 Results on constants redefinition
An overview of response accuracy is presented in Table 6, where we consider the hardest redefinitions ($R_{a}3$ and $R_{s}2$ for assignment and swapping respectively), as well as all three question levels, together with FF and MC response formats. It is observable that all tested LLMs, regardless of their size or model family, are prone to anchoring. This is especially evident in the FF format (since MC introduces a random choice factor), where models like Titan Large generate 60% anchored responses, while Claude Opus and Command r producing 47% and 53% respectively in this format.
@@
Claude sonnet v1 26.67 13.33 0.0 13.33 13.33 33.33 33.33 40.0 20.0 20.0 60.0 73.33
Claude sonnet v2 26.67 13.33 20.0 0.0 46.67 40.0 13.33 40.0 33.33 20.0 40.0 66.67
@@
To investigate the possible source of this phenomenon, we calculate the correlation between NR and post-redefinition anchored responses per LLM. Averaged results for all LLMs are presented in Table 7, revealing an intriguing pattern: for $Q_{1}$ and $Q_{2}$ levels, the correlation is either weak or negative. A negative correlation indicates that LLMs performing well in NR cases tend to anchor less. This suggests that when reasoning is of easy or medium level, LLMs are less likely to adhere rigidly to their default knowledge. This serves as a sanity check, confirming that LLMs understand the redefinition task and that anchoring rates are not due to prompting deficiencies. However, this trend reverses in the most challenging $Q_{3}$ level, particularly in the swapping cases. In these instances, a strong positive correlation is evident, implying that LLMs that originally perform well (thus are potent reasoners) tend to disregard the redefinition prompt and respond as they would in the NR case. This striking observation suggests that more capable reasoners anchor more on their prior knowledge. This pattern holds across both FF and MC formats, as well as different prompt types (more results in App. B).
@@
Anchoring is not only related to the per LLM reasoning capabilities, but also to the parameter size of the LLM itself. Even though larger models achieve higher correct response rate across redefinition levels, staying on par with their reasoning capabilities in the NR case, the anchoring rate also rises as LLM size increases. This indicates that larger models struggle to redefine well-known concepts, and instead rely on their existing knowledge. This is evident from Table 8, which shows the number of correct responses in the NR case versus the anchoring rate. For example, in the case of Llama, the 405B model anchors significantly when solving $Q_{3}$ questions post-redefinition compared to the smaller Llama70B, suggesting that the larger variant is less capable as a reasoner. The same trend holds for Mixtral8x7B and Mistral Large (123B): for the latter, anchoring is even higher compared to correct NR responses in the $Q_{3}$ level, meaning that the LLM provides the originally correct answer in the redefined problem (when this response is incorrect) more frequently than in the NR case (when the answer is correct). This unexpected behavior of Llama is further investigated under the MC response format for different prompting methods, focusing on the hardest redefinition ($R_{a}3$, $R_{s}2$) and question levels ($Q_{3}$). As illustrated in Figure 2, the anchoring rate rises with model scale, verifying the inverse scaling trend.
@@
Figure 4 presents results from two LLM families of known parameter count, Mistral and Llama, with varying sizes, for all redefinition levels on $Q_{3}$ questions, for FF and MC formats. The MC format is associated with higher anchoring rates (e.g., 73.33% and 93.33% for Llama 70B and 405B respectively) compared to FF responses (33.33% and 53.33%). This is rather expected, since the LLM is exposed to the default value of a constant, creating a conflict between memorization and instruction. The high probability the default value holds triggers the LLM to anchor to it, something that is not applicable in the FF case.
@@
Furthermore, prompting techniques play a crucial role in refusal rates, with FS mitigating refusal the most. This result is intuitive, as the LLM is exposed to more examples containing redefinitions in its input, making it less likely to refuse the task. Additional results are provided in App. D.
@@ 4.2 Results on units of measure redefinition
The findings on the redefinition of units of measure align with those of constants, also revealing a strong inverse trend: larger models (e.g., Mistral Large, Llama405B) consistently exhibit a higher anchoring rate compared to their smaller counterparts (Mistral7B, Llama8B), regardless of the response format, redefinition level, or prompting method. This trend is particularly notable in $Q_{3}$ questions. However, anchoring is relatively lower than in the constants case, likely because the actual values of units are not commonly used in calculations, making the LLMs less prone to anchoring. An thorough analysis is presented in App. E.
@@ Appendix B Model Knowledgeability
Claude sonnet v1 100.0 26.67 93.33 0.0 86.67 13.33 100.0 33.33 93.33 20.0 86.67 60.0
Claude sonnet v2 73.33 26.67 60.0 20.0 40.0 46.67 73.33 13.33 60.0 33.33 40.0 40.0
@@ Appendix C Inverse Trends
Figure 7 shows the rate of correct answers in the NR task, as well as the rates of correct, wrong, and anchored responses for LLaMA-8B and LLaMA-405B after the constant redefinition task in the MC response format. From this figure, it is observable that the same pattern as in Mistral (Figure 3) emerges, where the number of anchored responses is significantly higher in the larger model. Additionally, Figure 6 shows the number of anchored responses after constants redefinitions for models of varying sizes in the Mistral family in the MC response format. The trend is the same as in Llama models illustrated in the main paper, where larger variants tend to anchor more to their prior knowledge in comparison to the smaller ones, for both $R_{s}1,R_{s}2$ redefinition levels and all prompting techniques. Mixtral7x8B in the ZS prompting setup produces only slightly more anchored responses compared to the larger variant (Mistral Large (123B parameters)). However, this is likely due to the increased performance of the LLM in the NR case, as shown in Table 8. The difference is relatively small, meaning that while there is a measurable increase in anchoring for Mixtral7x8B, it is not substantial enough to indicate a significant deviation from the expected trend.
@@ Appendix D Refusal Rates
An important observation stated in the main paper is that completely wrong responses include instances where the LLM actively refuses to respond to the redefined problem, as analyzed before.
An analysis of the most interesting cases regarding wrong responses is presented in Figures 8(b), 9(b), 10(b), 11(b) for selected LLMs. For example, in the case of Llama8B (Figure 8(b)) the refusal rate diminishes as questions get harder. This reveals an overconfidence in responding instead of abstaining, leading to a problematic behavior, since Llama8B achieves very few correct responses in all question levels, and especially in harder ones ($Q_{3}$), as indicated in Figure 7(a). the difference between FF and MC response formats lies in the elevated number of blank responses in the MC case. This is because Llama8B suffers when prompted to select one of the predefined options, indicating high uncertainty. Nevertheless, when prompted to respond without restrictions, empty responses are almost non-existent, revealing another sense of overconfidence tied to response format.
Conversely, Llama70B (Figure 9(b)) refrains from empty responses, especially in harder cases, revealing its lower uncertainty in generating a response. Contrary to its smaller counterpart, its refusal rates decrease as questions get harder, leading to an increased number of wrong responses (reaching up to 100% in some cases) over refusals.
Finally, mixed patterns also occur in the case of Claude Sonnet v2 (Figure 11(b)), where alternating patterns between 100% refusal or 100% wrong responses are revealed (as in the $Q_{1}$ question level). Therefore, this model is rather unpredictable in whether it prefers to deny the task or respond erroneously, since there is no obvious reason behind this diverging behavior. Contrary to the aforementioned Mixtral8x7 case, Claude Sonnet v2 presents more wrong responses in the MC case in comparison to FF responses. That means that for some unexpected reason, it refuses to answer within an unrestricted setting, but results in wrong responses when presented with a limited number of options. This behavior reveals that Claude Sonnet v2 confidently performs redefinitions with ease and without much resistance, but fails to solve the redefined problem overall.
@@
Figure 9: Completely wrong responses breakdown for Llama70B. Blue denotes actually wrong responses, Purple indicates refusals, while Gray instances correspond to blank responses
@@
Figure 10: Completely wrong responses breakdown for Mixtral8x7. Blue denotes actually wrong responses, Purple indicates refusals, while Gray instances correspond to blank responses
(a) Response breakdown for Claude Sonnet v2 FF responses.
(b) Response breakdown for Claude Sonnet v2 MC responses.
Figure 11: Completely wrong responses breakdown for Claude Sonnet v2. Blue denotes actually wrong responses, Purple indicates refusals, while Gray instances correspond to blank responses
@@
The results indicate that models such as Command r+, Command light, and Claude Haiku consistently tend to provide responses, ignoring their validity, therefore exhibiting no refusal instances. This overconfidence is problematic in practice, even though useful in our experimentation, since reasoning shortcomings are exposed. In contrast, models such as Llama, Mistral, and Claude sonnet v2 occasionally decline to generate a response when faced with the redefined task, possibly acknowledging their intrinsic inability to answer properly. This observation showcases that those LLMs that prefer to abstain from responding are more robust, since a redefinition prompt could act as a malicious adversarial attack that aims to mislead the LLM towards generating invalid responses. On the other hand, this deliberate action prevents them to reveal their reasoning capabilities, once again verifying the trade-off between robustness and evident reasoning.
@@
Claude sonnet v2 ZS 0.0 0.0 25.0 100.0 16.67 0.0 0.0 0.0 0.0 0.0
CoT 0.0 0.0 14.29 100.0 0.0 0.0 100.0 0.0 0.0 50.0
FS 0.0 0.0 0.0 100.0 0.0 0.0 0.0 0.0 0.0 33.33
Table 14: The refusal rate for each LLM, method, and type of constants redefinitions for $Q_{1}$ questions.
@@
Table 15: The refusal rate for each LLM, method, and type of constants redefinition for $Q_{2}$ questions.
@@
Claude sonnet v2 ZS 14.29 16.67 16.67 50.0 25.0 0.0 0.0 16.67 33.33 0.0
CoT 0.0 14.29 14.29 20.0 25.0 0.0 0.0 0.0 0.0 0.0
FS 0.0 0.0 12.5 11.11 10.0 0.0 0.0 14.29 0.0 0.0
Table 16: The refusal rate for each LLM, method, and type of constants redefinition for $Q_{3}$ questions.
@@
Mistral7B 33.33 63.64 33.33 31.25 28.57 40.0 47.36 53.33 14.37 36.84 26.67 23.08
Mixtral8x7B 50.0 33.33 26.67 26.67 23.81 36.36 52.18 46.67 52.63 53.33 46.67 73.33
Mistral Large 33.33 20.0 26.67 26.67 60.37 66.67 100.0 53.33 52.24 40.0 73.33 66.67
Llama8B 0.0 52.18 0.0 42.11 25.28 52.94 71.43 40.9 42.11 50.0 28.2 31.43
Llama70B 8.2 13.33 0.0 0.0 25.71 40.0 33.33 46.67 15.21 46.67 38.46 73.33
Llama405B 0.0 0.0 0.0 13.33 26.67 53.33 26.67 46.67 6.67 20.0 58.82 93.33
Command text 13.33 20.0 6.67 6.67 6.67 26.67 40.0 26.67 13.33 26.67 13.33 42.85
Claude v2 32.66 13.33 20.0 0.0 51.22 44.45 15.58 40.0 33.33 22.22 47.06 66.67

### v2

@@ Abstract
Elena Stringli, Maria Lymperaiou, Giorgos Filandrianos, Athanasios Voulodimos, Giorgos Stamou School of Electrical and Computer Engineering, AILS Laboratory National Technical University of Athens ele191200@gmail.com, {marialymp, geofila}@islab.ntua.gr, thanosv@mail.ntua.gr, gstam@cs.ntua.gr
@@ 1 Introduction
Interestingly, inverse scaling is widely underexplored in literature. In this paper, we address this gap by examining the redefinition task, where well-known concepts are assigned alternative values, and LLMs are prompted to respond accordingly. For example, redefining $\pi=100$ (Figure 1), overwriting the default $\pi=3.14159$ refutes LLM’s prior knowledge, calling for flexible reasoning pathways in order to handle mathematical operations over the redefined $\pi$ value. Through vast experimentation on several redefinitions, LLM families and model scales, we conclude that:
@@ 2 Related work
is a fundamental concern, questioning the real barrier between LLMs and human cognition. While LLMs excel in linguistic competence, this ability is dissociated with thought Mahowald et al. (2024). In practice, LLMs are prone to performance degradation under alternative formulations, denoting their limited reasoning flexibility Wu et al. (2024) and their susceptibility to knowledge conflicts between contextual prompt information and stored facts Xu et al. (2024). Similar findings are reported in causal Jin et al. (2024); Gendron et al. (2024), analogical Lewis and Mitchell (2024); Stevenson et al. (2024) and commonsense Nezhurina et al. (2024) reasoning, where LLM performance declines sharply under diverging formulations. Alternative prompts are also shown to influence LLM capacity in arithmetic reasoning Ball et al. (2024); Li et al. (2024), translation over artificial languages and deductions with twists Li et al. (2024). Quite often, memorization accounts for reasoning, perplexing the evaluation of the real LLM abilities Xie et al. (2024); Lou et al. (2024); Wang et al. (2024a).
@@ 1 B bit (b) $8b$ $10b$ $6×10^{8}b$ $-4b$
We also design three levels of question difficulty. The first level ($Q_{1}$) mainly regards the question What is the first -non-zero- digit of {constant}?. The correct answer $A_{Q_{1}}$ is actually isolating the leftmost digit (ignoring leading zeros or the minus sign in cases of negative numbers) of the constant. For example, when $\pi$ has undergone the redefinition $\pi=500$ the correct response $A_{Q_{1}}$ is 5. There are some exceptions to the first digit $Q_{1}$, (presented in App. A). The next question level ($Q_{2}$), asks for a simple mathematical operation (e.g. What is $\pi$ multiplied by 3?), as presented in Table 2. The LLM has to execute this operation correctly to derive the correct $A_{Q_{2}}$, while the ground truth solution can be reached by utilizing a scientific calculator and the appropriate constant value. Finally, in the last and most difficult level ($Q_{3}$), questions requiring multi-hop reasoning are designed (e.g. What is the Earth’s surface area?), as the ones of Table 3.
@@
incorporates the following fundamental physical quantities: time (minutes-min), weight (kilogram-kg), length (meter-m) and light-year (ly), temperature (Kelvin-K), volume (milliliter-mL), energy (calorie-cal), pressure (atmosphere-atm), voltage (Volt-V), frequency (megaHz-MHz), force (newton-N), magnetic flux density (Tesla-T), area (hectare-ha), illuminance (lux-lx), and information (byte-B). We intervene on the scaling between each of those units and their derived counterparts for the same physical quantity: for example, a minute has 60 seconds, therefore a unit redefinition can be "redefine minutes to have 100 seconds". Details about such redefinitions are presented in Table 4.
As in the constants’ case, we offer three levels of questions difficulty. The easiest $Q_{1}$ level queries the actual conversion rule as defined in Physics, with a small adjustment to avoid the trivial case, where the answer lies in the prompt: instead of questioning How many seconds a minute is?, since its actual rephrasing exists in the prompt ("redefine a minute to have 100 seconds"), we prefer questions such as How many seconds are in two minutes?, imposing an undemanding calculation. In the $Q_{2}$ case, the LLM is tasked to solve an easy problem, applying fundamental physics equations or a unit scaling given minimal context. In the hardest $Q_{3}$ level, questions require more mathematical reasoning steps. All questions are illustrated in Table 5.
@@
MHz How many Hz are in 2 MHz? An oscillator operates at 4 MHz. What is the period of the wave in sec? A circuit has a signal with a frequency of 6 MHz. What is the wavelength of the signal if the speed of light is approximately $3\times 10^{8}$ m/s?
@@ 4 Experiments
We test 19 LLMs, including state-of-the-art (SoTA) model families: Llama 3 (8/70/405B), Mistral7B/Large/Mixtral8$\times$7b, Anthropic Claude (Opus/Instant/Haiku/v2/Sonnet 3.5&3.7), Cohere command (light/text/r/r+) and Amazon Titan (text lite/text express/large). All LLMs are prompted using zero shot (ZS), few shot (FS) and Chain of Thought (CoT) techniques.
@@
4. Completely wrong responses: The LLM produces blank, incorrect or inconsistent responses that do not fit any of the above cases. In some cases, it completely refuses to perform the redefinition.
To measure the impact of redefinitions, results post-redefinition are compared with those where no redefinition is performed (denoted as NR). We then focus on anchored responses, since they are mostly tied to the memorization versus reasoning trade-off in LLMs.
@@ 4.1 Results on constants redefinition
An overview of response accuracy is presented in Table 6, where we consider the hardest redefinitions ($R_{a}3$ and $R_{s}2$ for assignment and swapping respectively), as well as all three question levels, together with FF and MC response formats. It is observable that all tested LLMs, regardless of their size or model family, are prone to anchoring. This is especially evident in the FF format (since MC introduces a random choice factor), where models such as Titan Large generate 60% anchored responses, while Claude Opus and Command r produce 47% and 53% respectively in this format.
@@
Claude v2 26.67 13.33 20.0 0.0 46.67 40.0 13.33 40.0 33.33 20.0 40.0 66.67
Claude 3.5 Sonnet 26.67 13.33 0.0 13.33 13.33 33.33 33.33 40.0 20.0 20.0 60.0 73.33
Claude 3.7 Sonnet [^1] Without thinking module enabled for fair comparison. 0.0 0.0 0.0 6.67 13.33 13.33 33.33 20.0 6.67 20.0 40.0 33.33
@@
To investigate the possible source of this phenomenon, we calculate the correlation between NR rate and post-redefinition anchored responses per LLM. Averaged results for all LLMs are presented in Table 7, revealing an intriguing pattern: for $Q_{1}$ and $Q_{2}$ levels, the correlation is either weak or negative. A negative correlation indicates that LLMs performing well in NR cases tend to anchor less. This suggests that when reasoning is of easy or medium level, LLMs are less likely to adhere rigidly to their default knowledge. This serves as a sanity check, confirming that LLMs understand the redefinition task and that anchoring rates are not due to prompting deficiencies. However, this trend reverses in the most challenging $Q_{3}$ level, particularly in the swapping cases. In these instances, a strong positive correlation is evident, implying that LLMs that originally perform well (thus are potent reasoners) tend to disregard the redefinition prompt and respond as they would in the NR case. This striking observation suggests that more capable reasoners anchor more on their prior knowledge. This pattern holds across both FF and MC formats, as well as different prompt types. More results are provided in Appendix B.
@@
Anchoring is not only related to the per LLM reasoning capabilities, but also to the parameter size of the LLM itself. Even though larger models achieve higher correct response rate across redefinition levels, staying on par with their reasoning capabilities in the NR case, the anchoring rate also rises as LLM size increases. This indicates that larger models struggle to redefine well-known concepts, and instead rely on their existing knowledge. This is evident from Table 8, which shows the number of correct responses in the NR case versus the anchoring rate post-redefinition. For example, in the case of Llama, the 405B model anchors significantly when solving $Q_{3}$ questions post-redefinition compared to the smaller Llama70B, suggesting that the larger variant is less capable as a reasoner. The same trend holds for Mixtral8x7B and Mistral Large (123B): for the latter, anchoring is even higher compared to correct NR responses in the $Q_{3}$ level, meaning that the LLM provides the originally correct answer in the redefined problem (when this response is incorrect) more frequently than in the NR case (when the answer is correct). This unexpected behavior of Llama is further investigated under the MC response format for different prompting methods, focusing on the hardest redefinition ($R_{a}3$, $R_{s}2$) and question levels ($Q_{3}$). As illustrated in Figure 2, the anchoring rate rises with model scale, verifying the inverse scaling trend.
@@
Figure 4 presents results from two LLM families of known parameter count, Mistral and Llama, with varying sizes, for all redefinition levels on $Q_{3}$ questions, for FF and MC formats. The MC format is associated with higher anchoring rates (e.g., 73.33% and 93.33% for Llama 70B and 405B respectively) compared to FF responses (33.33% and 53.33%). This is rather expected, since the LLM is exposed to the default value of a constant in the presence of the correct MC candidate, creating a conflict between memorization and instruction. The high probability the default value holds triggers the LLM to anchor to it, something that is not applicable in the FF case, in which the LLM has to generate a response without any reference to the original value in the prompt.
@@
Furthermore, prompting techniques play a crucial role in refusal rates, with FS mitigating refusal the most. This result is intuitive, as the LLM is exposed to more examples containing redefinitions in its input, making it less likely to refuse the task. Additional results are provided in Appendix D.
(a) Claude 3.7 Sonnet without Thinking in the FF response format.
(b) Claude 3.7 Sonnet without Thinking in the MC response format.
(c) Claude 3.7 Sonnet with Thinking in the FF response format.
(d) Claude 3.7 Sonnet with Thinking in the MC response format.
Figure 6: Claude 3.7 Sonnet results without and with Thinking.
The impact of thinking
In the emergence of extended thinking modes in SoTA LLMs, such as Claude 3.7 Sonnet, we are able to elicit deeper reasoning processes of the model at hand, enabling better responses in multi-step reasoning situations, as in the case of elevated redefinition levels and question difficulty. Related findings are presented in Figure 6. As observed, the contribution of thinking is minimal, only slightly reducing the anchoring response rate in specific cases. Therefore, we can conclude that even advanced reasoning mechanisms cannot assist Claude 3.7 Sonnet in following the redefined thought processes needed to accurately respond, further strengthening the claim that the redefinition task challenges inherent reasoning functions and limitations of SoTA LLMs.
@@ 4.2 Results on units of measure redefinition
The findings on the redefinition of units of measure align with those of constants, also revealing a strong inverse trend: larger models (e.g., Mistral Large, Llama405B) consistently exhibit a higher anchoring rate compared to their smaller counterparts (Mistral7B, Llama8B), regardless of the response format, redefinition level, or prompting method. This trend is particularly notable in $Q_{3}$ questions. However, anchoring is relatively lower than in the constants case, likely because the actual values of units are not commonly used in calculations, making the LLMs less prone to anchoring. A thorough analysis is presented in Appendix E.
@@ 5 Conclusion
In terms of prompting, more refined strategies can be tested, especially in the FS case, where sophisticated prompting techniques have been recently proposed to enhance exemplar selection quality via reasoning similarity Panagiotopoulos et al. (2025) in place of the generic semantic similarity measure.
@@ Appendix B Model Knowledgeability
Claude v2 73.33 26.67 60.0 20.0 40.0 46.67 73.33 13.33 60.0 33.33 40.0 40.0
Claude 3.5 Sonnet 100.0 26.67 93.33 0.0 86.67 13.33 100.0 33.33 93.33 20.0 86.67 60.0
Claude 3.7 Sonnet 93.33 0.0 86.67 0.0 93.33 13.33 93.33 33.33 86.67 6.67 93.33 40.0
@@ Appendix C Inverse Trends
Figure 8 shows the rate of correct answers in the NR task, as well as the rates of correct, wrong, and anchored responses for Llama 8B and Llama 405B after the constant redefinition task in the MC response format. From this figure, it is observable that the same pattern as in Mistral (Figure 3) emerges, where the number of anchored responses is significantly higher in the larger model. Additionally, Figure 7 shows the number of anchored responses after constants redefinitions for models of varying sizes in the Mistral family in the MC response format. The trend is the same as in Llama models illustrated in the main paper (Figure 2), where larger variants tend to anchor more to their prior knowledge in comparison to the smaller ones, for both $R_{s}2,R_{a}3$ redefinition levels and all prompting techniques. Mixtral7x8B in the ZS prompting setup produces only slightly more anchored responses compared to the larger variant (Mistral Large (123B parameters)). However, this is likely due to the increased performance of the LLM in the NR case, as shown in Table 8. The difference is relatively small, meaning that while there is a measurable increase in anchoring for Mixtral7x8B, it is not substantial enough to indicate a significant deviation from the expected trend.
@@ Appendix D Refusal Rates
An important observation stated in the main paper is that completely wrong responses include instances where the LLM actively refuses to respond to the redefined problem.
An analysis of the most interesting cases regarding wrong responses is presented in Figures 9(b), 10(b), 11(b), 12(b) for selected LLMs. For example, in the case of Llama 8B (Figure 9(b)) the refusal rate diminishes as questions get harder. This reveals an overconfidence in responding instead of abstaining, leading to a problematic behavior, since Llama 8B achieves very few correct responses in all question levels, and especially in harder ones ($Q_{3}$), as indicated in Figure 8(a). The difference between FF and MC response formats lies in the elevated number of blank responses in the MC case. This is because Llama 8B suffers when prompted to select one of the predefined options, indicating high uncertainty. Nevertheless, when prompted to respond without restrictions, empty responses are almost non-existent, revealing another sense of overconfidence tied to response format.
@@
Figure 10: Completely wrong responses breakdown for Llama 70B. Blue denotes actually wrong responses, Purple indicates refusals, while Gray instances correspond to blank responses.
Conversely, Llama 70B (Figure 10(b)) refrains from empty responses, especially in harder cases, revealing its lower uncertainty in generating a response. Contrary to its smaller counterpart, its refusal rates decrease as questions get harder, leading to an increased number of actually wrong responses (reaching up to 100% in some cases) over refusals.
Finally, mixed patterns also occur in the case of Claude v2 (Figure 12(b)), where alternating patterns between 100% refusal or 100% actually wrong responses are revealed (as in the $Q_{1}$ question level). Therefore, this model is rather unpredictable in whether it prefers to deny the task or respond erroneously, since there is no obvious reason behind this diverging behavior. Contrary to the aforementioned Mixtral8x7 case, Claude v2 presents more wrong responses in the MC case in comparison to FF responses. That means that for some unexpected reason, it refuses to answer within an unrestricted setting, but results in wrong responses when presented with a limited number of options. This behavior reveals that Claude v2 confidently performs redefinitions with ease and without much resistance, but fails to solve the redefined problem overall.
@@
Figure 11: Completely wrong responses breakdown for Mixtral8x7. Blue denotes actually wrong responses, Purple indicates refusals, while Gray instances correspond to blank responses.
(a) Response breakdown for Claude v2 FF responses.
(b) Response breakdown for Claude v2 MC responses.
Figure 12: Completely wrong responses breakdown for Claude v2. Blue denotes actually wrong responses, Purple indicates refusals, while Gray instances correspond to blank responses.
@@
The results indicate that models such as Command r+, Command light, and Claude Haiku consistently tend to provide responses, ignoring their validity, therefore exhibiting no refusal instances. This overconfidence is problematic in practice, even though useful in our experimentation, since reasoning shortcomings are exposed. In contrast, models such as Llama, Mistral, and Claude v2 occasionally decline to generate a response when faced with the redefined task, possibly acknowledging their intrinsic inability to answer properly. This observation showcases that those LLMs that prefer to abstain from responding are more robust, since a redefinition prompt could act as a malicious adversarial attack that aims to mislead the LLM towards generating invalid responses. On the other hand, this deliberate action prevents them to reveal their reasoning capabilities, once again verifying the trade-off between robustness and evident reasoning.
@@
Claude 3.7 Sonnet ZS 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
CoT 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
FS 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Table 14: The refusal rate for each LLM, prompting technique, and type of constants redefinitions for $Q_{1}$ questions.
@@
Table 15: The refusal rate for each LLM, prompting technique, and type of constants redefinition for $Q_{2}$ questions.
@@
Claude 3.7 Sonnet ZS 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
CoT 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
FS 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Table 16: The refusal rate for each LLM, prompting technique, and type of constants redefinition for $Q_{3}$ questions.
@@
Mistral7B 33.33 50.0 33.33 28.57 28.57 40.0 45.45 53.33 14.28 35.71 26.67 23.08
Mixtral8x7B 38.46 33.33 26.67 26.67 23.08 35.71 36.37 46.67 46.15 53.33 46.67 73.33
Mistral Large 33.33 20.0 26.67 26.67 57.14 66.67 71.43 53.33 50.0 40.0 73.33 66.67
Llama8B 0.0 44.45 0.0 36.37 22.22 50.0 50.0 24.99 40.0 50.0 27.27 30.0
Llama70B 7.15 13.33 0.0 0.0 20.0 40.0 33.33 46.67 14.28 46.67 35.71 73.33
Llama405B 0.0 0.0 0.0 13.33 26.67 53.33 26.67 46.67 6.67 20.0 57.14 93.33
Command text 13.33 20.0 6.67 6.67 6.67 26.67 40.0 26.67 13.33 26.67 13.33 38.46
Claude v2 28.57 13.33 20.0 0.0 50.0 42.86 14.28 40.0 33.33 21.43 46.15 66.67